# answer ——–

### v1

@@ Abstract
We consider an alternative model based instead on the bootstrapped return, which adds to the partial return an estimate of the future return. Benefits of bootstrapped return over partial return and cumulative advantage models stem from their treatment of human beliefs. Unlike partial return, choices based on bootstrapped return reflect human beliefs about the environment. Further, while recovering the reward function from choices based on cumulative advantage requires that those beliefs are correct, doing so from choices based on bootstrapped return does not. In this sense, the bootstrapped return model disentangles goals from beliefs.
@@ 1 Introduction
To align an AI agent with human goals, it is common to design the agent to accumulate rewards that express human preferences. Manual specification of an effective reward function, however, is notoriously difficult (Akrour et al., 2012; Hadfield-Menell et al., 2016). And mispecification engenders reward hacking. Alignment can be improved via a reward function learned from human choice data (Akrour et al., 2012; Sadigh et al., 2017; Christiano et al., 2017; Ibarz et al., 2018; Brown and Niekum, 2019).
To generate choice data, it is common to present a human with pairs of action-observation trajectories and ask which they prefer. Learning from this data requires a model of human choice behavior. In existing models, the probability of choosing one trajectory over another is determined by scores assigned to the two, with each score calculated using a reward function. Christiano et al. (2017) popularized the use of partial return, expressed in the first row of Table 1. We use the term partial to indicate that this is the return over a partial trajectory, which is of finite length, as opposed to an infinite trajectory.
@@
Choices based on partial return are not influenced by human beliefs about the environment. This counters common sense. For example, consider the grid environment of Figure 1. Suppose the human attributes large reward to reaching the treasure and a small cost per timestep en route. Trajectory B incurs twice the cost of trajectory A; the partial returns are $-2$ for B and $-1$ for A. In spite of that, it is natural for a human to prefer whichever trajectory maximizes progress toward the treasure. If the human believes that the grid wraps around – so that after exiting to the left the agent appears on the right – then A is preferred. This is because the human’s quality assessment depends on beliefs about how each trajectory will subsequently play out. It is also in line with empirical results of (Christiano et al., 2017, Section 3.1.1), which demonstrate that learned reward functions express anticipated rewards.
To better reflect human choice behavior, Knox et al. (2024b) proposed the cumulative advantage [^1] Knox et al. (2024b) actually proposed regret, which is the negation of cumulative advantage. We use the latter so that we can more conveniently compare to other positive score formulas., as expressed in the second row of Table 1. Each advantage $\Delta_{*}(s_{t},a_{t})=Q_{*}(s_{t},a_{t})-V_{*}(s_{t})$ measures the amount by which the optimal value $Q_{*}(s_{t},a_{t})$ of the executed action exceeds the optimal state value $V_{*}(s_{t})$.
@@
To elucidate on how cumulative advantage relies on correct beliefs, let us revisit Figure 1. Suppose the agent is currently positioned as in the figure. Whether trajectory A or B is preferred depends on whether the human believes that the grid wraps around. Let us assume that the grid does not wrap around, but that the human believes it does and therefore prefers trajectory $A$. Then, an agent with correct knowledge of the environment will infer from advantages a large reward to the left.
@@
Let us revisit the example of Figure 1 to illustrate these benefits over partial return and cumulative advantage. While the partial return would encourage choice of A over B, either choice could be preferred according to the bootstrapped return. This is because depending on the human’s beliefs, the value estimate could be either large or small at the left end of the grid. In this way, the bootstrapped return accounts for the influence of human beliefs on choice.
Recall that when learning via the cumulative advantage model the agent can impute erroneous large reward at the left end of the grid. This happens if the human believes the grid wraps around but the agent knows it does not and assumes the human knows that as well. Suppose now that these beliefs are expressed through a bootstrapped reward model. Will the agent infer the correct reward function unlike with cumulative advantage? The answer is yes. In this case, there will be a large value estimate at the left end to reflect proximity to the treasure. And fitting a bootstrapped model to corresponding choice data recovers this value function and the human’s reward function $r$. In this sense, the bootstrapped return model disentangles goals, as expressed by $r$, from beliefs, which are expressed through the value function.
@@ 2.1 Trajectories and Lotteries
A lottery over infinite trajectories is a probability distribution. For example, for any policy $\pi$, transition matrix $P$, and partial trajectory $h$, the probability distribution $\mathbb{P}_{\pi,P}(\cdot|h)$ is a lottery. Given a lottery $\ell$ and function $f:\mathcal{H}_{\infty}\rightarrow\Re$, we denote by $\mathbb{E}_{\ell}[f(H_{\infty})]$ the expectation of $f(H_{\infty})$ with $H_{\infty}=(S_{0},A_{0},S_{1},A_{1},\ldots)$ sampled from $\ell$. Similarly, we denote by $\mathbb{E}_{\pi,P}[f(H_{\infty})|h]$ the expectation with respect to the lottery $\mathbb{P}_{\pi,P}(\cdot|h)$.
@@ 2.2 Preferences between Lotteries
Let $\succeq$ be a binary relation on $\mathcal{L}_{\infty}$. We interpret this relation as indicating the human’s preferences between infinite trajectory lotteries. For each infinite trajectory lottery $\ell\in\mathcal{H}_{\infty}$, discount factor $\gamma\in(0,1)$, and function $r:\mathcal{S}\times\mathcal{A}\rightarrow\Re$, let
@@
Loosely speaking, when $r$ expresses $\succeq$, lotteries with higher expected cumulative reward are preferred.
@@ 5 Closing Remarks
We have proposed and motivated a model of choice between partial trajectories based on bootstrapped return. Unlike with partial return, choices based on bootstrapped return are influenced by the human’s beliefs, and thus provide a more realistic model of human choice. When choices are made based on bootstrapped return, we show that it is still possible for the agent to infer the human’s goals, regardless of the human’s beliefs. In this sense, the agent can disentangle the human’s goals from their beliefs. This is unlike when choices are based on cumulative advantage, which relies on the human having correct beliefs about the environment.
@@
Our work adds to the growing chorus of work on RLHF emphasizing that current methods use unrealistic models of human behavior (see e.g., (Tien et al., 2023; Knox et al., 2024b) and (Casper et al., 2023) for a survey). Unlike previous works, our work focuses on how human beliefs about future rewards may influence choice between trajectories. One issue pertaining to human beliefs, but not addressed in our work, is that of partial observability which was recently discussed by Lang et al. (2024): when making comparisons between trajectories, the human often does not have full access to the agent’s trajectory. This may incentivize the agent to pursue trajectories such that the observation sequence revealed to the human looks good, while the underlying state sequence is actually not. Combining their work with a choice model based on bootstrapped return may be an interesting avenue for future work.
@@ Acknowledgements
This research was supported by a grant from the US Army Research Office. We thank Derya Cansever for stimulating conversations and encouragement on this work.
@@ A.2 Relative Value
is well-defined and finite. Further, $\tilde{v}_{\ell,r}$ does not depend on $h$.
@@ A.3 Lottery Preferences via Relative Value
If $r$ expresses $\succeq$ then, for all $\ell,\ell^{\prime}\in\mathcal{L}_{\infty}$,
@@
$\tilde{v}_{\ell,r}-\tilde{v}_{\ell,r}=$ $\lim_{\gamma\uparrow 1}\left(\mathbb{E}_{\pi,P}\left[\sum_{t=0}^{\infty}\gamma^{t}\tilde{r}_{\ell,r}(S_{t},A_{t})\Big{|}h\right]-\mathbb{E}_{\pi,P}\left[\sum_{t=0}^{\infty}\gamma^{t}\tilde{r}_{\ell^{\prime},r}(S_{t},A_{t})\Big{|}h\right]\right)$
$=$ $\lim_{\gamma\uparrow 1}\left(\mathbb{E}_{\pi,P}\left[\sum_{t=0}^{\infty}\gamma^{t}(r_{\ell,r}(S_{t},A_{t})-\overline{r}_{\ell,r})\Big{|}h\right]-\mathbb{E}_{\pi,P}\left[\sum_{t=0}^{\infty}\gamma^{t}(r_{\ell^{\prime},r}(S_{t},A_{t})-\overline{r}_{\ell,r})\Big{|}h\right]\right)$
$=$ $\lim_{\gamma\uparrow 1}\left(\mathbb{E}_{\pi,P}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{\ell,r}(S_{t},A_{t})\Big{|}h\right]-\mathbb{E}_{\pi,P}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{\ell^{\prime},r}(S_{t},A_{t})\Big{|}h\right]\right)$
@@
Hence, $\tilde{v}_{\ell,r}\geq\tilde{v}_{\ell,r}$ if and only if $\lim_{\gamma\uparrow 1}(v_{\ell,r,\gamma}-v_{\ell^{\prime},r,\gamma})\geq 0$. ∎
@@ A.5 The Alignment Theorem
We first establish Statement 1. By Lemma 5, for all lotteries $\ell$ and $\ell^{\prime}$ induced by some $(\pi,P)\in\mathcal{P}$ with $h=(s_{0},a_{0},\ldots,s_{T})$ and $h^{\prime}=(s^{\prime}_{0},a^{\prime}_{0},\ldots,s^{\prime}_{T^{\prime}})$, respectively, $\ell\succeq\ell^{\prime}$ if and only if $\tilde{v}_{\ell,r_{*}}\geq\tilde{v}_{\ell^{\prime},r_{*}}$. Lemma 4 then implies that $\sum_{t=0}^{T-1}\tilde{r}_{\pi,P,r_{*}}(s_{t},a_{t})+\tilde{V}_{\pi,P,r_{*}}(s_{T})\geq\sum_{t=0}^{T^{\prime}-1}\tilde{r}_{\pi,P,r_{*}}(s^{\prime}_{t},a^{\prime}_{t})-\tilde{V}_{\pi,P,r_{*}}(s^{\prime}_{T^{\prime}})$. It follows that $(\tilde{r}_{\pi,P,r_{*}},\tilde{V}_{\pi,P,r_{*}})$ expresses $\succeq_{\partial}$.

### v3

@@ Abstract
We consider an alternative model based instead on the bootstrapped return, which adds to the partial return an estimate of the future return. Benefits of bootstrapped return over partial return and cumulative advantage models stem from their treatment of human beliefs. Unlike partial return, choices based on cumulative advantage or bootstrapped return reflect human beliefs about the environment. Further, while recovering the reward function from choices based on cumulative advantage requires that those beliefs are correct, doing so from choices based on bootstrapped return does not. In this sense, the bootstrapped return model disentangles the human’s goals from their beliefs.
@@ 1 Introduction
To align an AI agent with human goals, it is common to design the agent to accumulate rewards that express human preferences. Manual specification of an effective reward function, however, is notoriously difficult (Akrour et al., 2012; Hadfield-Menell et al., 2016). And mispecification leads to reward hacking. Alignment can be improved via a reward function learned from human choice data (Akrour et al., 2012; Sadigh et al., 2017; Christiano et al., 2017; Ibarz et al., 2018; Brown and Niekum, 2019).
To generate choice data, it is common to present a human with pairs of action-observation trajectories and ask which they prefer. Inferring the human’s goals from this data requires a model of human choice behavior. In existing models, the probability of choosing one trajectory over another is determined by scores assigned to the two, with each score calculated using a reward function that expresses the human’s goals. Christiano et al. (2017) popularized the use of partial return, expressed in the first row of Table 1. We use the term partial to indicate that this is the return over a partial trajectory, which is of finite length, as opposed to an infinite trajectory.
@@
Choices based on partial return are not influenced by human beliefs about the environment. This counters common sense. For example, consider the grid environment of Figure 1. Suppose the human’s goal is for the robot to reach the treasure as quickly as possible. We can model this goal as the human attributing large reward to reaching the treasure and a small cost per timestep en route. Then, trajectory B incurs twice the cost of trajectory A; the partial returns are $-2$ for B and $-1$ for A. In spite of that, it is natural for a human to prefer whichever trajectory maximizes progress toward the treasure. Under reasonable beliefs about the environment, this could be trajectory B. However, if the human believes that the grid wraps around – so that after exiting to the left the agent appears on the right – then A is preferred. This is because the human’s quality assessment depends on beliefs about how each trajectory will subsequently play out. As another example, consider a human choosing between chess trajectories. Assume that the human’s sole goal is winning. It seems reasonable to suppose that the human’s evaluation of a trajectory will be influenced by how the human thinks the game will subsequently play out, given the chess board at the end of the trajectory. More generally, it seems plausible that human choices are influenced by what they believe will be the continuation of a partial trajectory. This is in line with empirical results of (Christiano et al., 2017, Section 3.1.1), which demonstrate that when using the partial return model the learned reward functions express anticipated rewards. This suggests that, in practice, choices are made not only based on partial return, but also on anticipated future rewards.
To better reflect human choice behavior, Knox et al. (2024b) proposed the cumulative advantage [^1] Knox et al. (2024b) actually proposed regret, which is the negation of cumulative advantage. We use the latter so that we can more conveniently compare to other positive score formulas., as expressed in the second row of Table 1. Each advantage $\Delta_{*}(s_{t},a_{t})=Q_{*}(s_{t},a_{t})-V_{*}(s_{t})$ measures the the difference between the optimal value $Q_{*}(s_{t},a_{t})$ of the executed action and the optimal state value $V_{*}(s_{t})$.
@@
To elucidate on how cumulative advantage relies on correct beliefs, let us revisit Figure 1. Suppose the agent is currently positioned as in the figure. Whether trajectory A or B is preferred depends on whether the human believes that the grid wraps around. Let us assume that the grid does not wrap around, but that the human believes it does and therefore prefers trajectory $A$. Then, an agent with correct knowledge of the environment will infer from advantages a large reward to the left. Thus, the agent will infer a reward function that does not express the human’s goals.
@@
Let us revisit the example of Figure 1 to illustrate these benefits over partial return and cumulative advantage. While the partial return would encourage choice of A over B, either choice could be preferred according to the bootstrapped return. This is because depending on the human’s beliefs, the value estimate could be either large or small at the left end of the grid. As another example, consider a human choosing between chess trajectories. Under the bootstrapped return model, the human’s evaluation of a trajectory will be influenced by how the human thinks the game will subsequently play out, given the chess board at the end of the trajectory. In this way, the bootstrapped return accounts for the influence of human beliefs on choice.
Recall that when learning via the cumulative advantage model the agent can impute erroneous large reward at the left end of the grid. This happens if the human believes the grid wraps around but the agent knows it does not and assumes the human knows that as well. Suppose now that these beliefs are expressed through a bootstrapped reward model. Will the agent infer the correct reward function unlike with cumulative advantage? The answer is yes. In this case, there will be a large value estimate at the left end to reflect proximity to the treasure. And fitting a bootstrapped model to corresponding choice data recovers this value function and the human’s reward function $r$. In this sense, the bootstrapped return model disentangles goals, as expressed by $r$, from beliefs, which are expressed through the value function. This is crucial. For instance, in the game of chess, while the human may often choose trajectories in which the opponent’s queen is captured, the agent ought to learn a reward function that reflects that the human’s sole goal is to win. The human only wishes that the queen is captured in so far as the human believes that it will lead to winning down the line.
@@ 2.1 Trajectories and Lotteries
A lottery over infinite trajectories is a probability distribution over infinite trajectories. For example, for any policy $\pi$, transition matrix $P$, and partial trajectory $h$, the probability distribution $\mathbb{P}_{\pi,P}(\cdot|h)$ is a lottery. Given a lottery $\ell$ and function $f:\mathcal{H}_{\infty}\rightarrow\Re$, we denote by $\mathbb{E}_{\ell}[f(H_{\infty})]$ the expectation of $f(H_{\infty})$ with $H_{\infty}=(S_{0},A_{0},S_{1},A_{1},\ldots)$ sampled from $\ell$. Similarly, we denote by $\mathbb{E}_{\pi,P}[f(H_{\infty})|h]$ the expectation with respect to the lottery $\mathbb{P}_{\pi,P}(\cdot|h)$.
@@ 2.2 Preferences between Lotteries
Let $\succeq$ be a binary relation on $\mathcal{L}_{\infty}$. We interpret this relation as indicating the human’s preferences between infinite trajectory lotteries. For each infinite trajectory lottery $\ell\in\mathcal{L}_{\infty}$, discount factor $\gamma\in(0,1)$, and function $r:\mathcal{S}\times\mathcal{A}\rightarrow\Re$, let
@@
Loosely speaking, when $r$ expresses $\succeq$, lotteries with higher expected cumulative reward are preferred. The above preference structure corresponds to comparing lotteries based on a so-called bias-optimality criterion Puterman (2014, See Section 5.4.3). If one lottery has a higher average reward, that lottery is preferred. If the two lotteries have equal average reward, then the lottery with the highest so-called bias is preferred.
@@ 5 Closing Remarks
We have proposed and motivated a model of choice between partial trajectories based on bootstrapped return. Unlike with partial return, choices based on bootstrapped return are influenced by the human’s beliefs, and thus provide a more realistic model of human choice. When choices are made based on bootstrapped return, we show that it is still possible for the agent to infer the human’s goals, regardless of the human’s beliefs. In this sense, the agent can disentangle the human’s goals (as expressed by the reward function) from their beliefs about future observations. This is unlike when choices are based on cumulative advantage, which relies on the human having correct beliefs about the environment.
@@
Our work adds to the growing chorus of work on RLHF emphasizing that current methods use unrealistic models of human behavior (see, e.g., (Tien et al., 2023; Knox et al., 2024b) and, for a survey, (Casper et al., 2023)). Unlike previous works, our work focuses on how human beliefs about future rewards may influence choice between trajectories. One issue pertaining to human beliefs, but not addressed in our work, is that of partial observability which was recently discussed by Lang et al. (2024): when making comparisons between trajectories, the human often does not have full access to the agent’s trajectory. This may incentivize the agent to pursue trajectories such that the observation sequence revealed to the human looks good, while the underlying state sequence is actually not. Combining their work with a choice model based on bootstrapped return may be an interesting avenue for future work.
There is a clear connection between our work and Savage (1972)’s axiomatization of subjective expected utility maximization. In his framework, goals are represented by a utility function rather than a reward function, and beliefs by a subjective probability distribution over states, rather than over future infinite trajectories. While the choice setting considered by Savage is very different from ours, it shares the basic idea of choices being influenced by both goals and beliefs. One interesting avenue for future work is to, in Savage’s spirit, derive our two axioms in terms of more basic axioms that reference neither reward nor probability.
@@ Acknowledgements
This research was supported by a seed fund from the Stanford Institute for Human-Centered AI, contributed by SCBX and SCB10X, and a grant from the US Army Research Office. We thank Derya Cansever, Micah Carroll, Stephane Hatgis-Kessell, Anmol Kagrecha, Saurabh Kumar, Ramesh Johari, and Tanwa Arpornthip for stimulating conversations and helpful feedback.
@@ A.2 Relative Value
is well-defined and finite.
@@ A.3 Lottery Preferences via Relative Value
If $r$ expresses $\succeq$ then, for all lotteries $\ell$ and $\ell^{\prime}$ induced by some $(\pi,P)\in\mathcal{P}$ with $h=(s_{0},a_{0},\ldots,s_{T})$ and $h^{\prime}=(s^{\prime}_{0},a^{\prime}_{0},\ldots,s^{\prime}_{T^{\prime}})$, respectively,
@@
$\tilde{v}_{\ell,r}-\tilde{v}_{\ell^{\prime},r}=$ $\lim_{\gamma\uparrow 1}\left(\mathbb{E}_{\pi,P}\left[\sum_{t=0}^{\infty}\gamma^{t}\tilde{r}_{\ell,r}(S_{t},A_{t})\Big{|}h\right]-\mathbb{E}_{\pi,P}\left[\sum_{t=0}^{\infty}\gamma^{t}\tilde{r}_{\ell^{\prime},r}(S_{t},A_{t})\Big{|}h\right]\right)$
$=$ $\lim_{\gamma\uparrow 1}\left(\mathbb{E}_{\pi,P}\left[\sum_{t=0}^{\infty}\gamma^{t}(r_{\ell,r}(S_{t},A_{t})-\overline{r}_{\ell,r})\Big{|}h\right]-\mathbb{E}_{\pi,P}\left[\sum_{t=0}^{\infty}\gamma^{t}(r_{\ell^{\prime},r}(S_{t},A_{t})-\overline{r}_{\ell^{\prime},r})\Big{|}h\right]\right)$
=$\stackrel{{\scriptstyle(a)}}{{=}}$ $\lim_{\gamma\uparrow 1}\left(\mathbb{E}_{\pi,P}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{\ell,r}(S_{t},A_{t})\Big{|}h\right]-\mathbb{E}_{\pi,P}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{\ell^{\prime},r}(S_{t},A_{t})\Big{|}h\right]\right)$
@@
Because $(\pi,P)\in\mathcal{P}$, we have that $\overline{r}_{\ell,r}=\overline{r}_{\ell^{\prime},r}$, which means that $(a)$ follows. Since $\tilde{v}_{\ell,r}-\tilde{v}_{\ell^{\prime},r}=\lim_{\gamma\uparrow 1}(v_{\ell,r,\gamma}-v_{\ell^{\prime},r,\gamma})$, we have that $\tilde{v}_{\ell,r}\geq\tilde{v}_{\ell^{\prime},r}$ if and only if $\lim_{\gamma\uparrow 1}(v_{\ell,r,\gamma}-v_{\ell^{\prime},r,\gamma})\geq 0$. ∎
@@ A.5 The Alignment Theorem
We first establish Statement 1. By Axiom 2 there exists $(\pi,P)\in\mathcal{P}$ such that for all partial trajectories $h,h^{\prime}$,
$h\succeq_{\partial}h^{\prime}\qquad$ $\Leftrightarrow\qquad\underbrace{\mathbb{P}_{\pi,P}(\cdot|h)}_{\ell}\succeq\underbrace{\mathbb{P}_{\pi,P}(\cdot|h^{\prime})}_{\ell^{\prime}}$
$\stackrel{{\scriptstyle(a)}}{{\Leftrightarrow}}\qquad\tilde{v}_{\ell,r_{*}}\geq\tilde{v}_{\ell^{\prime},r_{*}}$
$\stackrel{{\scriptstyle(b)}}{{\Leftrightarrow}}\qquad\sum_{t=0}^{T-1}\tilde{r}_{\pi,P,r_{*}}(s_{t},a_{t})+\tilde{V}_{\pi,P,r_{*}}(s_{T})\geq\sum_{t=0}^{T^{\prime}-1}\tilde{r}_{\pi,P,r_{*}}(s^{\prime}_{t},a^{\prime}_{t})-\tilde{V}_{\pi,P,r_{*}}(s^{\prime}_{T^{\prime}})$
The steps (a) and (b) follows by Lemma 5, and Lemma 4, respectively. It follows that $(\tilde{r}_{\pi,P,r_{*}},\tilde{V}_{\pi,P,r_{*}})$ expresses $\succeq_{\partial}$. Note that we use $\ell$ and $\ell^{\prime}$ as shorthand for the lotteries induced by $(\pi,P)\in\mathcal{P}$ with $h$ and $h^{\prime}$, respectively.
@@ A.6 Relationships between Advantages and Rewards
Appendix B Measure-Theoretic Considerations
In the main text, we left out details on exactly what sigma-algebra the lotteries are defined on. Further, in the main text we used expectations with respect to those lotteries, but never discussed whether these expectations are well-defined. In this section, we fill in the details, and show that the expectations we use in the paper are well-defined.
B.1 Measurable Space
To formally define probability distributions over infinite trajectories, we need to define a measurable space that we can define our measures on. To do so, we first equip $\mathcal{H}_{\infty}$ with the product topology denoted by $\mathcal{T}(\mathcal{H}_{\infty})$. We then define our sigma-algebra to be the Borel-sigma with respect to $\mathcal{T}(\mathcal{H}_{\infty})$. We denote this sigma-algebra by $\mathcal{B}(\mathcal{H}_{\infty})$. All probability distributions are defined with respect to the measurable space $\left(\mathcal{H}_{\infty},\mathcal{B}(\mathcal{H}_{\infty})\right)$. Note that for all partial trajectories $h_{t}\in\mathcal{H}$ the prefix set $\{h_{\infty}^{\prime}\in\mathcal{H}_{\infty}|h_{t}^{\prime}=h_{t}\}$ is measurable.
B.2 Expectations are well-defined
In the paper, we compute expectations $\mathbb{E}_{\ell}[g(H_{\infty})]$ with respect to some lottery $\ell$ and function $g:\mathcal{H}_{\infty}\to\Re$. For instance, we often compute this expectation with $g=\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})$, for some $\gamma\in(0,1)$ and $r:\mathcal{S}\times\mathcal{A}\to\Re$. We will prove that for this choice of $g$ this expectation is well-defined.
We will define an infinite sequence of functions $\{g_{T}\}$ as follows. For all $T$, let $g_{T}:\mathcal{H}_{\infty}\to\Re$ be defined as $g_{T}(h_{\infty})=\sum_{t=0}^{T-1}\gamma^{t}r(s_{t},a_{t})$ for all $h_{\infty}\in\mathcal{H}_{\infty}$. Then,
  1. 1.
    $g_{T}$ converges to $g$ pointwise as $T\to\infty$.
  2. 2.
    $g_{T}$ is absolutely bounded by $\sum_{t=0}^{\infty}\gamma^{t}\max_{s,a}|r(s,a)|$.
  3. 3.
    $g_{T}$ is measurable.
Therefore, by the Dominated Convergence Theorem, $g$ is integrable, and the expectation is thus well-defined. The same argument can be made for when $g$ computes the average reward or the state visitation frequency.